# Nash Equilibria in Games with Playerwise Concave Coupling Constraints: Existence and Computation

### v2

@@ Abstract
We study the existence and computation of Nash equilibria in continuous static games where the players’ admissible strategies are subject to shared coupling constraints, i.e., constraints that depend on their joint strategies. Specifically, we focus on a class of games characterized by playerwise concave utilities and playerwise concave constraints. Prior results on the existence of Nash equilibria are not applicable to this class, as they rely on strong assumptions such as joint convexity of the feasible set. By leveraging topological fixed point theory and novel structural insights into the contractibility of feasible sets under playerwise concave constraints, we give an existence proof for Nash equilibria under weaker conditions. Having established existence, we then focus on the computation of Nash equilibria via independent gradient methods under the additional assumption that the utilities admit a potential function. To account for the possibly nonconvex feasible region, we employ a log barrier regularized gradient ascent with adaptive stepsizes. Starting from an initial feasible strategy profile and under exact gradient feedback, the proposed method converges to an $\epsilon$-approximate constrained Nash equilibrium within $\mathcal{O}(\epsilon^{-3})$ iterations.
@@ 1 INTRODUCTION
Optimization problems involving multiple self-interested agents with coupled objectives frequently arise in applications across robotics, computer science, and economics. While in standard game theoretic frameworks, the players’ feasible action sets are decoupled, in many interactive tasks, players are restricted to optimizing over a set of admissible strategies that are influenced by the choices of others. This game setting is captured by introducing coupling constraints, that is, constraints which depend on the players’ joint strategy profile. For instance, in economic modeling, such constraint games have been applied to analyze equilibria under resource limitations and market constraints (Arrow and Debreu,, 1954). In other applications, players may be subject to joint safety requirements such as collision avoidance in multi-robot control (Gu et al.,, 2023), or operating within a shared budget, such as limited bandwidth in communication networks (Lasaulce and Tembine,, 2011). Despite their high relevance, several questions about the existence and computation of equilibria in this class of games remain open. In this paper, we address some of these questions, focusing on concave games with playerwise concave coupling constraints. To clarify our contributions, we first discuss the state of the art on games with coupling constraints.
@@
Shortly after Nash’s foundational work on existence of mixed strategy equilibria in noncooperative finite action games, Debreu (Debreu,, 1952) introduced games with coupling constraints. In his formulation, each player may face a coupling constraint distinct from those of other players—a setting we refer to as non-shared coupling constraints. The equilibrium of such a game, called a social equilibrium and later referred to as a constrained Nash equilibrium, is one where no player has an incentive to unilaterally deviate within her set of feasible actions. Debreu showed that such equilibria exist in games with compact, convex strategy spaces and playerwise quasi-concave utilities, provided that each player’s set of feasible responses remains nonempty, irrespective of the strategies chosen by others. This requirement is often restrictive: in many practical scenarios, such as under shared safety constraints, violations caused by a subset of players may leave others unable to restore feasibility. However, this assumption appears necessary in Debreu’s setting, as without it one can construct examples—even with linear utilities and non-shared linear coupling constraints—where no constrained Nash equilibrium exists (see Appendix B).
Later, Rosen, (1965) showed that for the class of constrained games in which all players have concave utilities and share the same coupling constraints, constrained Nash equilibria exist without requiring the nonemptiness of feasible response sets as in Debreu, (1952). However, this work assumed a jointly convex feasible region. This assumption is restrictive, since even bilinear coupling constraints in two-player settings are generally nonconvex. Such a bilinear structure arises naturally for instance in mixed strategy extensions of finite action games. Thus, an open question is whether existence of Nash equilibria can be guaranteed in games with shared coupling constraints under conditions weaker than joint convexity or nonemptiness of feasible response sets?
Playerwise concavity—a relaxation of joint concavity—is a common structural assumption on utilities that ensures the existence of Nash equilibria in games without constraints (Rosen,, 1965). We consider the natural extension of this game class to settings with coupling constraints and prove existence of constrained Nash equilibria for games with playerwise concave utilities and playerwise concave coupling constraints.
@@
While existence results are fundamental, for Nash equilibria to be plausible outcomes of multi-agent interactions, they should also be computationally tractable. It is known that adding constraints to games can introduce significant computational challenges. For instance, it was recently shown that convex-concave zero-sum two-player games with coupling constraints are computationally intractable (Daskalakis et al.,, 2021; Bernasconi et al.,, 2024), a stark contrast to their unconstrained counterparts.
Similar to prior existence results, much of the literature on computing constrained Nash equilibria relies on restrictive assumptions about the game structure. For example, in monotone games with a jointly convex feasible region, coupling constraints can be handled using the variational inequality framework (Facchinei and Pang,, 2003). These assumptions on both utilities and constraints are stronger than playerwise concavity.
Unfortunately, even for shared bilinear constraints, finding a feasible strategy is NP-hard (Witsenhausen,, 1986). However, when an initial feasible strategy is given, one can ask whether there exist learning dynamics—understood here as iterative update procedures—under which players converge to a constrained Nash equilibrium. Under this initial feasibility assumption, recent work has provided positive answers in the context of constrained Markov potential games (Alatur et al.,, 2024; Jordan et al.,, 2024), a setting with finite action spaces. We provide such a convergent learning method for the class of continuous action potential games with playerwise concave utilities and shared playerwise concave coupling constraints.
@@
    For the class of games with playerwise convex feasible regions, we establish the existence of constrained Nash equilibria without requiring nonemptiness of feasible response sets. Our approach generalizes the result of Rosen, (1965) by relaxing the assumption on the feasible region from joint convexity to playerwise convexity.
    Insights. We build on tools from topological fixed point theory, specifically the works of Eilenberg and Montgomery, (1946), and Begle, (1950), which extend fixed point results from convex to more general, contractible domains. A key technical contribution of our work is the analysis of the topological structure of the feasible region, where we show that playerwise concavity of constraints yields a contractibility property.
@@
    Restricting the above class to potential games with coupling constraints, we develop an independent learning algorithm with provable convergence to a constrained Nash equilibrium at an $\mathcal{O}(\epsilon^{-3})$ rate. Players simultaneously update their strategies using only gradient feedback on their own utility and the shared constraints, without requiring additional communication.
    Insights. We draw inspiration from log barrier regularized interior point methods. Originally proposed by Frisch, (1955), this approach has been extensively studied in the context of constrained optimization problems (Megiddo,, 1989; Hinder and Ye,, 2024), but has remained unexplored in the context of games. Starting from an initial feasible strategy, and by carefully selecting adaptive learning rates, we guarantee convergence along the potential function to approximate equilibria.
@@ 2 PRELIMINARIES
In this paper, we study static games with continuous strategy spaces. Let $[m]$ be the set of players. For each $i\in[m]$, $\mathcal{X}_{i}\subset\mathbb{R}^{d_{i}}$ is a compact convex set of available strategies. Joint strategy profiles are denoted by $x=(x_{1},\dots,x_{m})\in\mathcal{X}\coloneqq\times_{i\in[m]}\mathcal{X}_{i}$. Upon playing strategy $x\in\mathcal{X}$, each player $i$ receives utility $u_{i}(x)$ where $u_{i}:\mathcal{X}\to\mathbb{R}$ is continuous. Shared coupling constraints are introduced through a set of functions and thresholds $\{(c_{j},\alpha_{j})\}_{j\in[b]}$ where for each $j\in[b]$, $\alpha_{j}\in\mathbb{R}$, and $c_{j}:\mathcal{X}\to\mathbb{R}$ is a continuous function. Players are assumed to be utility maximisers subject to lower bounds on the constraint functions. The set of strategy profiles considered feasible is given by
@@
For joint strategies of all players other than $i$, we use the notation $x_{-i}\coloneqq(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{m})\in\mathcal{X}_{-i}\coloneqq\times_{j\in[m]\setminus\{i\}}\mathcal{X}_{j}$. For $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, let $\mathcal{C}_{i}(x_{-i})\coloneqq\{x_{i}\in\mathcal{X}_{i}\mid(x_{i},x_{-i})\in\mathcal{C}\}$ denote the feasible response set of player $i$.
Nash equilibria
For games without coupling constraints, a Nash equilibrium is a strategy profile $x\in\mathcal{X}$ such that for all $i\in[m]$ and playerwise deviations $x^{\prime}_{i}\in\mathcal{X}_{i}$, we have $u_{i}(x^{\prime}_{i},x_{-i})\leq u_{i}(x)$. In our constrained setting, $x\in\mathcal{X}$ is called constrained Nash equilibrium if $x\in\mathcal{C}$ and for all $i\in[m]$ and $x^{\prime}_{i}\in\mathcal{C}_{i}(x_{-i})$, it holds that $u_{i}(x^{\prime}_{i},x_{-i})\leq u_{i}(x)$.
Playerwise concavity of utilities is commonly assumed for proving existence of unconstrained Nash equilibria. We additionally make an equivalent playerwise concavity assumption on the constraints.
Assumption 1.
For all $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, suppose $x_{i}\mapsto u_{i}(x_{i},x_{-i})$ and $x_{i}\mapsto c_{j}(x_{i},x_{-i})$ are concave for all $j\in[b]$.
@@ 3 EXISTENCE OF CONSTRAINED NASH EQUILIBRIA
In this section, we establish our existence result for constrained Nash equilibria. We begin by discussing the structure of the feasible region induced by playerwise concave constraints. We then identify minimal conditions that guarantee existence of equilibria and illustrate how these results generalize prior work. Finally, we give an overview of our proof approach.
@@ 3.1 Structure of the feasible region
Existence proofs for Nash equilibria typically proceed by applying a fixed point theorem to a continuous best response function. In order to define this best response map over players’ feasible response sets, for each player $i$, $\mathcal{C}_{i}(x_{-i})$ should vary continuously as a set-valued function of $x_{-i}$. The following example demonstrates that playerwise concavity of constraints is not sufficient for this continuity property to hold. Moreover, Example 1 shows that constrained Nash equilibria may not exist in games with feasible regions that lack such continuity of the feasible response map.
Figure 1: Illustration of the feasible sets $\mathcal{C}^{(1)}$ and $\mathcal{C}^{(2)}$ described in Example 1 and Example 2, respectively.
Example 1.
Consider a two-player game with continuous strategy spaces $\mathcal{X}_{1}=\mathcal{X}_{2}=[0,1]$ and feasible set $\mathcal{C}^{(1)}$ defined by a constraint $c_{1}(x_{1},x_{2})\coloneqq(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})$ with threshold $\alpha_{1}=0$; illustrated on the left-hand side of Figure 1. Moreover, define the players’ utilities as $u_{1}(x_{1},x_{2})=u_{2}(x_{1},x_{2})\coloneqq x_{1}(1-x_{2})$. Observe that no strategy $(x_{1},x_{2})\in\mathcal{C}^{(1)}$ with $x_{1}<\frac{1}{2}$ can be a constrained Nash equilibrium, since then for $x_{1}^{\prime}=\frac{1}{2}$, we have $u_{1}(x_{1}^{\prime},x_{2})>u_{1}(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2})\in\mathcal{C}^{(1)}$. By an analogous argument, no $(x_{1},x_{2})\in\mathcal{C}^{(1)}$ with $x_{2}>\frac{1}{2}$ can be a constrained Nash equilibrium. The only point not covered by above two cases is $(\frac{1}{2},\frac{1}{2})$ which however is not a Nash equilibrium either.
We note that in Example 1, the discontinuity in the set-valued mapping $x_{1}\mapsto\mathcal{C}_{2}(x_{1})$ occurs at $x_{1}=\frac{1}{2}$ where the map is upper but not lower-semicontinuous (see Appendix C for definitions). In order to rule out such degeneracies, we require that for all points on the boundary of the feasible region $\mathcal{C}$, there exists a uniform descent direction for all constraints. Together with continuity of the constraints $c_{j}(\cdot)$, this ensures that for any point on the boundary of $\mathcal{C}$ and any player $i$, there exists a neighborhood of feasible responses for player $i$ that is stable under small perturbations of the other players’ strategies.
Assumption 2.
@@
We point out that this condition is a playerwise variant of the Mangasarian-Fromovitz constraint qualification condition (Mangasarian and Fromovitz,, 1967).
Next, in order to demonstrate the generality of our conditions, we present an example of a constrained game whose feasible region $\mathcal{C}^{(2)}$ satisfies our Assumptions 1 and 2, but does not fulfill the joint convexity or nonempty feasible response requirements imposed by previous existence results. Feasible regions similar to the one in Example 2 naturally arise in matrix games, where nonconvexity stems from joint probability products, and the nonemptiness condition fails when satisfying the constraint requires some coordination among players to choose certain feasible joint actions.
Example 2.
Let again $\mathcal{X}_{1}=\mathcal{X}_{2}=[0,1]$ and consider the feasible set $\mathcal{C}^{(2)}$ defined by two constraints, $c_{1}(x_{1},x_{2})\coloneqq(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})$ and $c_{2}(x_{1},x_{2})\coloneqq-(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})$, together with thresholds fixed, for concreteness, at $\alpha_{1}=\frac{1}{111},\alpha_{2}=-\frac{1}{11}$, as shown on the right-hand side of Figure 2. The utility structure is irrelevant for this example. The resulting feasible set $\mathcal{C}^{(2)}$ is illustrated as the blue region on the right-hand side of Figure 1. The constraints satisfy Assumption 1, as fixing $x_{1}$ yields a linear function in $x_{2}$, and vice versa. Moreover, Assumption 2 holds, as $|\mathcal{J}(x)|=1$ for all $x\in\mathcal{C}^{(2)}$ and neither $c_{1}$ nor $c_{2}$ has any critical points on the boundary of $\mathcal{C}^{(2)}$. One can also observe that $\mathcal{C}^{(2)}$ is not convex, hence Rosen, (1965)’s result is not applicable. Moreover, for any $x\in\mathcal{X}$ with $x_{1}=0.5$, we have $\mathcal{C}^{(2)}_{2}(x_{1})=\emptyset$. Hence, results relying on the nonemptiness of feasible responses such as Debreu, (1952) are not applicable either.
In the following, we address the existence of Nash equilibria for games with feasible regions such as in Example 2, namely the class of constrained games captured by Assumptions 1 and 2.

### v3

@@ Abstract
We study the existence and computation of Nash equilibria in concave games where the players’ admissible strategies are subject to shared coupling constraints. Under playerwise concavity of constraints, we prove existence of Nash equilibria. Our proof leverages topological fixed point theory and novel structural insights into the contractibility of feasible sets, and relaxes strong assumptions for existence in prior work. Having established existence, we address the question of whether in the presence of coupling constraints, playerwise independent learning dynamics have convergence guarantees. We address this positively for the class of potential games by designing a convergent algorithm. To account for the possibly nonconvex feasible region, we employ a log barrier regularized gradient ascent with adaptive stepsizes. Starting from an initial feasible strategy profile and under exact gradient feedback, the proposed method converges to an $\epsilon$-approximate constrained Nash equilibrium within $\mathcal{O}(\epsilon^{-3})$ iterations.
Machine Learning, ICML
@@ 1 Introduction
Optimization problems involving multiple self-interested agents with coupled objectives frequently arise in applications across economics (Krawczyk & Uryasev, 2000), computer science (Menache & Ozdaglar, 2011; Sahiner et al., 2022), and robotics (Wang et al., 2019). While in standard game theoretic frameworks, the players’ feasible action sets are decoupled, in many interactive tasks, players are restricted to optimizing over a set of admissible strategies that are influenced by the choices of others. This game setting is captured by introducing coupling constraints, that is, constraints which depend on the players’ joint strategy profile. For instance, in economic modeling, such constraint games have been applied to analyze equilibria under resource limitations and market constraints (Arrow & Debreu, 1954). In other applications, players may be subject to joint safety requirements such as collision avoidance in multi-robot control (Gu et al., 2023), or operating within a shared budget, such as limited bandwidth in communication networks (Lasaulce & Tembine, 2011). In machine learning, convex-concave min-max problems with shared coupling constraints have recently appeared in the analysis of Generative Adversarial Networks (Sahiner et al., 2022).
Despite their high relevance, several questions about the existence and computation of equilibria in games with coupling constraints remain open. The central challenge is that coupling constraints destroy the separability of players’ feasible response sets, significantly complicating existence and computation of equilibria.
In this paper, we consider games with playerwise concave coupling constraints and playerwise concave utilities. We prove the existence of equilibria and design an algorithm that provably converges to an equilibrium. To place our contributions, we discuss the state-of-the-art on existence and computation of equilibria for the considered class of games.
@@
Shortly after Nash’s foundational work on the existence of mixed-strategy equilibria in finite games with decoupled action sets (Nash, 1950), Debreu introduced games with coupling constraints (Debreu, 1952). In his formulation, each player may face a distinct coupling constraint, a setting we refer to as non-shared coupling constraints. A constrained Nash equilibrium is then defined as a strategy profile from which no player has an incentive to unilaterally deviate within her feasible set.
For non-shared constraints, Debreu established existence of constrained Nash equilibria under playerwise quasi-concavity of utilities. However, his proof is based on the assumption that each player’s feasible set is nonempty regardless of the strategies chosen by others. This assumption is restrictive: in many practical scenarios, such as safety or threshold constraints on players’ strategies, violations induced by a subset of players may leave others with no feasible response.
For shared coupling constraints, where all players face the same constraints, Rosen showed that constrained Nash equilibria exist under playerwise concavity of utilities (Rosen, 1965). While this result did not require the restrictive assumption of Debreu (1952) on each player’s feasible response set, it relied on the assumption that the feasible region is jointly convex.
The joint convexity assumption limits the applicability of the result: even simple bilinear coupling constraints lead to non-jointly convex feasible regions.
The above results on existence of equilibria in games with coupling constraints raise the question of whether the equilibrium existence can be guaranteed for games with shared coupling constraints under weaker conditions on the feasible response sets or on the the joint strategy space.
We resolve this question for the class of games with playerwise concave constraints, a relaxation of joint concavity that allows constraints to depend on other players’ actions in a nonconcave manner. Playerwise concavity is a natural assumption: just as playerwise concavity of utilities suffices for equilibrium existence in unconstrained games, playerwise concavity of constraints ensures convexity of individual feasible sets without imposing restrictive joint structure.
@@
While existence results are fundamental for Nash equilibria to be plausible outcomes of multi-agent interactions, practical relevance also requires computational tractability. Without further assumptions on the game structure, computing Nash equilibria is intractable even in two-player games (Daskalakis et al., 2009; Papadimitriou et al., 2023). Consequently, positive algorithmic results exist for structured classes of games. In particular, in the absence of coupling constraints, the class of potential games (Heliou et al., 2017; Anagnostides et al., 2022) and zero-sum games (Mokhtari et al., 2020; Mertikopoulos et al., 2019) admit efficient algorithms and well-understood learning dynamics.
By contrast, in the presence of coupling constraints, the computational landscape remains far less understood, even under favorable structural assumptions such as in zero-sum or potential games.
It has been shown that convex-concave zero-sum two-player games with non-shared coupling constraints are computationally intractable (Daskalakis et al., 2021; Bernasconi et al., 2024); a stark contrast to their unconstrained counterparts. Tractability can be recovered when constraints are shared and the feasible region is jointly convex, where the problem can be address through well-studied variational inequality framework (Facchinei & Pang, 2003).
Beyond joint convexity, however, little is known. In fact, even when constraints are shared and feasible sets are playerwise rather than jointly convex, finding a feasible strategy is NP-hard (Witsenhausen, 1986). This rules out algorithmic approaches that rely on first solving feasibility and motivates a different perspective: Starting from an initial feasible strategy, can players’ learning dynamics converge to a constrained Nash equilibrium?
We pursue this direction for potential games with playerwise concave utilities and shared playerwise concave coupling constraints. In the absence of coupling constraints, independent learning dynamics is known to converge to Nash equilibria in such games (Anagnostides et al., 2022). We show that even in the presence of shared coupling constraints, there exists a playerwise decoupled learning algorithm that converges to an equilibrium.
Summarizing the above, our contributions are as follows.
@@
    For the class of games with playerwise convex feasible regions, we establish the existence of constrained Nash equilibria without requiring nonemptiness of feasible response sets (Debreu, 1952) or the joint convexity of the feasible region (Rosen, 1965).
    A key technical contribution of our work is a novel proof technique for analysing the topological structure of the feasible region. This allows us to build on tools from topological fixed point theory, specifically the works of Eilenberg & Montgomery (1946), and Begle (1950), which extend fixed point results from convex to more general, contractible domains.
@@
    Restricting the above class to potential games with coupling constraints, we develop a playerwise decoupled learning algorithm with provable convergence. In our approach, players simultaneously update their strategies using only gradient feedback on their own utility and the shared constraints, without requiring coordination or communication.
    Our algorithmic novelty is to extend a log barrier interior point method (Frisch, 1955) from single-player constrained optimization (Megiddo, 1989; Hinder & Ye, 2024) to the multi-player game setting. Starting from an initial feasible strategy, and by carefully selecting adaptive learning rates, we guarantee convergence to an $\epsilon$-approximate constrained Nash equilibrium at an $\mathcal{O}(\epsilon^{-3})$ rate.
@@ 2 Preliminaries
In this paper, we study static games with continuous strategy spaces. Let $[m]$ be the set of players. For each $i\in[m]$, $\mathcal{X}_{i}\subset\mathbb{R}^{d_{i}}$ is a compact convex set of available strategies. Joint strategy profiles are denoted by $x=(x_{1},\dots,x_{m})\in\mathcal{X}\coloneqq\times_{i\in[m]}\mathcal{X}_{i}$. Upon playing strategy $x\in\mathcal{X}$, each player $i$ receives utility $u_{i}(x)$ where $u_{i}:\mathcal{X}\to\mathbb{R}$ is continuous. Shared coupling constraints are introduced through a set of functions and thresholds $\{(c_{j},\alpha_{j})\}_{j\in[b]}$ where for each $j\in[b]$, $\alpha_{j}\in\mathbb{R}$, and $c_{j}:\mathcal{X}\to\mathbb{R}$ is a continuous function. Players are assumed to be utility maximizers subject to lower bounds on the constraint functions. The set of strategy profiles considered feasible is given by
@@
For joint strategies of all players other than $i$, we use the notation $x_{-i}\coloneqq(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{m})\in\mathcal{X}_{-i}\coloneqq\times_{j\in[m]\setminus\{i\}}\mathcal{X}_{j}$. For $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, let
$\mathcal{C}_{i}(x_{-i})\coloneqq\{x_{i}\in\mathcal{X}_{i}\mid(x_{i},x_{-i})\in\mathcal{C}\}$
denote the feasible response set of player $i$.
Definition 2.1.
A constrained Nash equilibrium is a feasible strategy $x\in\mathcal{C}$ such that for all $i\in[m]$ and $x^{\prime}_{i}\in\mathcal{C}_{i}(x_{-i})$, we have $u_{i}(x^{\prime}_{i},x_{-i})\leq u_{i}(x)$.
Playerwise concavity of utilities is commonly assumed for proving existence of Nash equilibria even without constraints. We naturally make an equivalent playerwise concavity assumption on the constraints.
Assumption 2.2.
For all $i\in[m]$, $x_{-i}\in\mathcal{X}_{-i}$, and $j\in[b]$, $x_{i}\mapsto u_{i}(x_{i},x_{-i})$ and $x_{i}\mapsto c_{j}(x_{i},x_{-i})$ are concave.
Remark
Restricting our attention to shared constraints is motivated by the fact that in the absence of additional assumptions—such as those imposed in Debreu (1952)—a constrained Nash equilibrium need not exist; see Appendix B for an example.
@@ 3 Existence of Constrained Nash Equilibria
In this section, we establish our existence result for constrained Nash equilibria. To this end, we introduce an additional assumption on the feasible set, which, for playerwise concave constraints, generalizes the existence conditions of prior work. Our key insight is that the feasible response set must have a well-behaved boundary (see Example 3.2); this can be achieved with conditions less stringent than prior work, as shown through Example 3.3.
@@ 3.1 Structure of the Feasible Region
We first introduce the following assumption, which ensures that at every boundary of the feasible region $\mathcal{C}$, there exists a uniform descent direction for all constraints.
Assumption 3.1.
@@
The above condition is a playerwise variant of the Mangasarian-Fromovitz constraint qualification condition (Mangasarian & Fromovitz, 1967), widely employed in single-player constrained optimization. The following example illustrates a case in which Assumption 3.1 does not hold.
Example 3.2.
Consider a two-player game with continuous strategy spaces $\mathcal{X}_{1}=\mathcal{X}_{2}=[0,1]$ and feasible set $\mathcal{C}^{(1)}$ defined by $c_{1}(x_{1},x_{2})\coloneqq(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})$ with threshold $\alpha_{1}=0$; illustrated on the left-hand side of Figure 1. Note that at the point $\hat{x}=(\frac{1}{2},\frac{1}{2})\in\mathcal{C}$, we have $\nabla_{x_{1}}c_{1}(\hat{x})=\nabla_{x_{2}}c_{1}(\hat{x})=0$, implying that Assumption 3.1 is violated.
Figure 1: Illustration of the feasible sets $\mathcal{C}^{(1)}$ and $\mathcal{C}^{(2)}$ described in Example 3.2 and Example 3.3, respectively.
The main challenge in Example 3.2 is that at $\hat{x}=(\frac{1}{2},\frac{1}{2})$, the set-valued feasible response mappings $x_{1}\mapsto\mathcal{C}_{2}(x_{1})$ and $x_{2}\mapsto\mathcal{C}_{1}(x_{2})$ are upper but not lower-semicontinuous (see Appendix C for definitions). This leads to discontinuity of the best-response maps, and as a result, fixed-point arguments for continuous self-maps no longer apply to guarantee existence of an equilibrium.
Next, in order to demonstrate the generality of our conditions, we present an example of a constrained game whose feasible region $\mathcal{C}^{(2)}$ satisfies both our Assumptions 2.2 and 3.1, but does not fulfill the joint convexity or nonempty feasible response requirements imposed by previous existence results. Non-jointly convex feasible regions of this type arise naturally in mixed strategy extensions of matrix games (as defined in Example 4.2).
Example 3.3.
Let again $\mathcal{X}_{1}=\mathcal{X}_{2}=[0,1]$ and consider the feasible set $\mathcal{C}^{(2)}$ defined by two constraints, $c_{1}(x_{1},x_{2})\coloneqq(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})$ and $c_{2}(x_{1},x_{2})\coloneqq-(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})$, together with thresholds fixed, for concreteness, at $\alpha_{1}=\frac{1}{111},\alpha_{2}=-\frac{1}{11}$, as shown on the right-hand side of Figure 3.3. The utility structure is irrelevant for this example. The resulting feasible set $\mathcal{C}^{(2)}$ is illustrated as the blue region on the right-hand side of Figure 1.
Observe that the constraints satisfy Assumption 2.2, as fixing $x_{1}$ yields a linear function in $x_{2}$, and vice versa. Moreover, Assumption 3.1 holds, as $|\mathcal{J}(x)|=1$ for all $x\in\mathcal{C}^{(2)}$ and neither $c_{1}$ nor $c_{2}$ has any critical points on the boundary of $\mathcal{C}^{(2)}$. Furthermore, note that $\mathcal{C}^{(2)}$ is not jointly convex, hence Rosen (1965)’s result is not applicable. Moreover, for any $x\in\mathcal{X}$ with $x_{1}=0.5$, we have $\smash{\mathcal{C}^{(2)}_{2}(x_{1})=\emptyset}$. Hence, results relying on the nonemptiness of feasible responses such as Debreu (1952) are not applicable either.
Having motivated and clarified our setting, we address the existence of Nash equilibria for the class of constrained games captured by Assumptions 2.2 and 3.1.